OpenStudy (anonymous):

26. The time t required to drive a certain distance varies with the speed r. If it takes 4 hours to drive the distance at 35 miles per hour, how long will it take to drive the same distance at 45 miles per hour? (1 point) about 3.11 hours 140 hours about 5.14 hours 393.75 hours

OpenStudy (anonymous):

3.11

OpenStudy (anonymous):

d=v*t 4*35=45*t sft

OpenStudy (anonymous):

thank you! :)